Simplify the expression:
√3𝑥^4∙√9𝑥/√3𝑥^3

Answer by Boreal(15235)

If this is sqrt(3x^4)*sqrt(9x)/sqrt(3x^3), then the answer is sqrt(27x^5/3x^3)=sqrt (9x^2)=3x